OpenStudy (anonymous):

q^2+3q-54=(q-6)(q+[ ]) FIND: [ ] n^2+5n-50=(n+5)(n-[ ]) FIND: [ ] please :) thanks!

OpenStudy (austinl):

\[q^2 +3q-54=(q-6)(q+[x])\] \[n^2 + 5n -50 = (n+5)(n-[x])\] In both instances you are factoring.

OpenStudy (austinl):

In the first one, how would you factor the left side?

OpenStudy (austinl):

What multiplied by 6 gives you negative 54, and when added gives you positive 3?

OpenStudy (austinl):

negative 6, excuse me

OpenStudy (anonymous):

9!

OpenStudy (anonymous):

ahh thank you

OpenStudy (austinl):

Right, sorry I got distracted. So, what would the answer to the second one be?

OpenStudy (anonymous):

-10 :)

OpenStudy (austinl):

Well, because there is a negative sign already there, it should just be x=10